# SpeechCap: Leveraging Playful Impact Captions to Facilitate Interpersonal Communication in Social Virtual Reality

### v1

@@ 7.2. Procedure
The user study was conducted individually for each participant. The study session for a participant was run in a one-on-one manner in which an experimenter (either the first or second author of this paper) serves the participant to go through the entire procedure. The study sessions lasted 75 to 90 minutes and were held in a lab space in our university. After the study, the participant received a gift card valued \$50 in local currency. For analysis, the study processes were fully audio-taped and the activities in VR were screen-recorded. Regarding hardware settings, two sets of Meta Quest 2 devices and two laptop computers (paired with Bluetooth microphones) were used for the experimenter and the participant in a session. An additional router was configured to host a local area network to serve the multi-user session of the SpeechCap’s VR application (subsubsection 5.1.4).
The core design of this user study is to evaluate SpeechCap through two typical conversation scenarios in social VR: (1) a private conversation between two close friends and (2) a public discussion on a debatable topic with multiple participants. First, the dyad conversation scenario simulates a private conversation in which rich emotional expressions occur. In this scenario, the experimenter and the participant play as close friends to talk about their daily life and personal feelings with the help of impact captions. In the multi-person conversation scenario, the participant will join an ongoing discussion to listen to others and then provide his/her own opinion. In the two scenarios, participants will have opportunities to play the role of both speaker (i.e., information sender) and listener (i.e., information receiver) in turns, allowing them to play with various impact caption instances in very different visual and interactive appearances. Specifically, a full user study session consists of four sections as follows:
@@ 7.3. Results and Findings of Post-study Survey
The post-study survey questionnaires consist of two sections. The first section employs four questions that ask whether the impact caption design space is understandable, interesting, and meaningful. The second section contains six questions about the user experience with SpeechCap, considering enjoyment, emotional expression, clarity of presentation, and overall experience of having conversations with impact captions in VR. The questions are designed in a 7-point Likert scale format  (Joshi et al., 2015), requiring the participants to respond with their level of agreement on the statements.
@@
In the first section (Figure 7 upper), the participants agreed that the overall design of impact captions to be used for communication in VR was understandable (Q1, M=6.36, SD=0.74) and interesting (Q2, M=6.14, SD=0.95). For the visual appearances and design of individual impact captions, 12 participants agreed that the integration of non-textual elements (ornaments, emojis, speech bubbles) with texts was meaningful (Q3, M=5.71, SD=1.07). And 13 participants believed that the visual design (i.e., color, typeface, and size) of textual elements was reasonable (Q4, M=5.79, SD=0.89). No participants responded with disagreement on the design of impact captions.
In the second section (Figure 7 lower), most of the participants (11/14) agreed that their experiences with SpeechCap were somewhat enjoyable (Q5, M=5.50, SD=1.22). From an information sender’s perspective, participants diverged on the effectiveness of using impact captions for emotional expression in social VR (Q6, M=4.29, SD=1.20). Six participants agreed that impact captions can help express their emotions effectively, while the other five committed objections. Two participants responded with a neutral opinion. This may relate to the ambiguity caused by the mapping rules for non-verbal visual elements. For the next question about whether impact captions were helpful for presentation clarity (Q7, M=4.43, SD=0.76), six participants responded with agreement, while five participants stayed neutral and one responded “somewhat disagreed”. From a receiver’s perspective, the majority of the participants (13/14) agreed that they could perceive the emotions of others by seeing and playing with the impact captions in VR (Q8, M=5.43, SD=0.65). Most of the participants (12/14) also agreed that impact captions could improve presentation clarity and help them understand other people in VR (Q9, M=5.36, SD=0.74). In addition, ten participants agreed that impact captions were useful in setting the tone of conversation (Q10, M=5.00, SD=0.96).
@@ 7.4. Results and Findings of Semi-structured Interviews
During the user study, all participants fully experienced the two scenarios and played the roles of both information sender and receiver in conversations. Creative usages of impact captions in VR emerged from their play-through. In the semi-structured interviews, we discussed questions regarding the impact captions’ design space and user experiences of SpeechCap with the participants based on their individual performance. Overall, the participants responded positively, indicating that the impact captions were “interesting and enjoyable to play with.” (P1, P2, P4, P10, P13). Two authors analyzed the transcribed scripts of the interviews by iterative open coding  (Corbin and Strauss, 2014) and finally revealed the following findings showing the meaningfulness of our impact captions’ design space, and the usefulness of SpeechCap system.
@@ 7.4.1. F1: Textual Elements Can Effectively Support Conversations in Social VR
Participants generally believed that visible texts in impact captions enhance their conversational experience in VR (P2, P3, P4, P6, P7, P11, P12, P13, P14). On one hand, text elements help users accurately express complex or profound meanings. P12 emphasized the expressive capability of impact captions: “The texts within impact captions have a high potential in expressiveness, because texts can convey complex, abstract concepts as well as accurately represent ironic or literary content, which non-textual elements cannot achieve.
@@
7.4.2. F2: Non-textual Elements are Engaging but Possibly Ambiguous
@@ 7.4.4. F4: Needs for Identifying Keywords in Conversations
Although the participants mostly provided us with positive feedback to SpeechCap, they also pointed out some drawbacks. One of the most commonly asked questions by the participants when they started using SpeechCap was about how we decided the words to be rendered as impact captions, as they usually found there were too many impact captions emerging when speaking continuously. After knowing our simple POS-tag-based filtering method to remove functional words, most participants suggested an improvement on it. P2 suggested, “Don not turn every word into an impact caption. Please extract the key information (of the speech) and only show impact captions of the relevant keywords.” Other participants, like P11, thought too much text content increased the reading and comprehension costs and then reduced the overall experience of having a conversation with SpeechCap. P11 complained, “No one wants to read that much text when using VR.” Particularly, in the multi-person conversation scenario, as more impact captions were spawned and accumulated in limited space, some participants (P1, P2) reported that they felt the field of vision was partially obstructed.
@@ 8. Discussion
8.1. Designing Visual Rhetoric with Interactivity for Facilitating Communication in Social VR
As suggested by previous research, mediating interpersonal communication in virtual spaces is actually about supporting people in constructing cognitive and affective connections  (McVeigh-Schultz and Isbister, 2021; Palmer, 1995) and discovering self-identity beyond the real world  (Freeman and Maloney, 2021; Sykownik et al., 2022; Maloney et al., 2020). Although most of the known communication methods in social VR are still imperfect  (Wei et al., 2024; Dzardanova et al., 2022; Akselrad et al., 2023; Sykownik et al., 2023; Tanenbaum et al., 2020; Baker et al., 2021).
@@ 8.1.1. Highlighting the Visual Rhetoric and Interactivity
To fill this gap, our work incorporates speech-driven captioning mechanisms into social VR with improvements on captions’ visual design and interactivity. Inspired by the original impact captions used in entertaining TV shows  (Sasamoto, 2014), we design and implement a form of visual rhetoric with an integration of textual and non-textual elements to convey verbal and non-verbal information simultaneously and provide several human-caption interactions to support speech conversations in social VR. The user study results confirmed that our impact-caption-mediated method for communication can effectively enhance clarity (subsubsection 7.4.1) and reinforce non-verbal cues (subsubsection 7.4.2) in social VR conversations. Moreover, we also found that the rich interactivity offered by impact captions can promote interpersonal interactions between users (subsubsection 7.4.3). Particularly, SpeechCap presents a concrete example of how expressive visual objects with interactivity can mediate communication and further support social interactions in social VR.
@@ 8.1.2. Considering Generalizability
As for the generalizability of our impact-caption-mediated method, the aforementioned application scenarios (section 6) demonstrate that SpeechCap can be widely applied to different cases to support user needs in social VR. It can enhance emotional expressions to create engaging and intimate atmospheres, support live presentations for teaching scientific knowledge in educational scenarios, and further provide a promising captioning mechanism that can bring immersive social VR experiences to deaf and hard-of-hearing (DHH) individuals.
8.1.3. Research Directions for Future Work
@@ 8.2. Understanding the Ambiguity in SpeechCap
However, the inevitable ambiguity is a double-edged sword in digital media design. While bringing risks of miscommunication, it can also facilitate creativity  (Stacey and Eckert, 2003), support negotiation  (Gaver et al., 2003), and even enhance the human-likeness of AI agents for human-AI communication  (Liu et al., 2024). Therefore, a trade-off arises: for CMC tools, designers must balance the clarity of communication while maintaining certain ambiguity to meet other design goals  (Stacey and Eckert, 2003; Aoki and Woodruff, 2005). This trade-off leads to two directions of dealing with the ambiguity: mitigating and utilizing.
@@ 8.3.1. Presenting Verbal Information Selectively
Verbal information is fundamental in interpersonal communication  (Liu et al., 2023). However, when enhancing verbal information for social VR conversations, we should consider selective presentations rather than making all the words in speech visible. In our user study, participants suggested avoiding presenting every word in speech, since too many insignificant words could be distracting and cause unexpected visual clutter (subsubsection 7.4.4). Although we have already designed a mechanism to softly control the number of impact captions, trying to minimize the possibility of visual clutter (subsubsection 5.3.4), the POS-tag-based filtering mechanism in the text processor of SpeechCap is too simple to be an effective keyword extraction method. With the fact evolution of AI and NLP technologies, future work can explore the ways to involve advanced NLP technologies such as Large Language Models (LLMs)  (Maragheh et al., 2023) to achieve keyword extraction with accurate understandings of the speech content and the user’s intentions.
@@ 8.3.2. Using Visual Rhetoric to Present Non-verbal Information
Non-verbal information plays an important role in communication, social interactions  (McVeigh-Schultz and Isbister, 2022; Maloney et al., 2020; Aburumman et al., 2022; Liebman and Gergle, 2016), and self-presentations  (Freeman and Maloney, 2021; Sykownik et al., 2022; Zhang et al., 2022) in social VR. Our work demonstrates that rich visual elements can effectively convey non-verbal information in social VR, showcasing the potential of visual rhetoric design for communication tools. However, we currently only use simple mapping rules to decide how captions correspond to specific visual elements in SpeechCap. These rules are inadequate for sufficiently and accurately expressing the various types of non-verbal information, as pointed out by participants in the user study. Therefore, future research can start from seeking in-depth understandings of the associations between specific visual elements and non-verbal information (such as color-emotion association  (Hanada, 2018; Wilms and Oberfeld, 2018)), and create novel visual rhetoric designs with effectiveness and elegance for supporting various applications of social VR.
@@ 9. Conclusion
This research demonstrates the potential of impact captions as a novel tool to facilitate interpersonal communication in social VR. By integrating both verbal and non-verbal cues through dynamic and playful typographic elements, impact captions enrich the interpersonal interactions in real-time conversations in virtual spaces. Our proof-of-concept system, SpeechCap, showcases the feasibility and benefits of this approach, while also identifying areas for improvement. Future work could focus on refining the design of impact captions to minimize ambiguity and exploring customization features to better meet diverse user needs and enhance inclusiveness. Overall, our impact-caption-inspired approach offers new avenues for engaging and meaningful interactions among social VR users, paving the way for more immersive and effective communication experiences.

### v3

@@ 7.2. Procedure
The user study was conducted in a one-on-one format, where an experimenter (either the first or second author of this paper) guided each participant through the procedure. The study sessions lasted 75 to 90 minutes and were held in a university lab space. After the study, the participant received a gift card valued \$50 in local currency. For analysis, the study processes were fully audio-taped and the activities in VR were screen-recorded. Each session utilized two Meta Quest 2 devices and two laptops, each paired with a Bluetooth microphone, for the experimenter and the participant. An additional router was configured to host a local area network to serve the multi-user session of SpeechCap’s VR application (subsubsection 5.1.4).
The primary goal of this user study was to evaluate SpeechCap using two common social VR scenarios: (1) a private conversation between two friends and (2) a public discussion on a debatable topic. The dyad conversation scenario simulated a private conversation designed to evoke rich emotional expressions. In this scenario, the experimenter and the participant play as close friends to talk about their daily life and personal feelings with the help of impact captions. In the multi-person conversation scenario, the participant joined an ongoing discussion, listened to others, and then shared their own opinion. In the two scenarios, participants will have opportunities to play the role of both speaker (i.e., information sender) and listener (i.e., information receiver) in turns, allowing them to play with various impact caption instances in very different visual and interactive appearances. Specifically, a full user study session consists of four sections as follows:
@@ 7.3. Results and Findings of Post-study Survey
The post-study survey questionnaires consisted of two sections. The first section employs four questions that ask whether the impact caption design space is understandable, interesting, and meaningful. The second section contains six questions about the user experience with SpeechCap, considering enjoyment, emotional expression, clarity of presentation, and overall experience of having conversations with impact captions in VR. The questions are designed in a 7-point Likert scale format  (Joshi et al., 2015), requiring the participants to respond with their level of agreement on the statements.
@@
In the first section (Figure 7 upper), participants agreed that the design of impact captions for VR communication was both understandable (Q1, M=6.36, SD=0.74) and interesting (Q2, M=6.14, SD=0.95). Regarding the visual design of individual captions, 12 participants agreed that combining non-textual elements (ornaments, emojis, speech bubbles) with text was meaningful (Q3, M=5.71, SD=1.07). And 13 participants believed that the visual design (i.e., color, typeface, and size) of textual elements was reasonable (Q4, M=5.79, SD=0.89). No participants responded with disagreement on the design of impact captions.
In the second section (Figure 7 lower), most of the participants (11/14) agreed that their experiences with SpeechCap were somewhat enjoyable (Q5, M=5.50, SD=1.22).
From an information sender’s perspective, participants diverged on the effectiveness of using impact captions for emotional expression in social VR (Q6, M=4.29, SD=1.20). Six participants agreed that impact captions can help express their emotions effectively, while the other five committed objections. Two participants responded with a neutral opinion. For the next question about clarity (Q7, M=4.43, SD=0.76), six participants responded with agreement, while five participants stayed neutral and one responded “somewhat disagreed”. The results of these two questions indicate that SpeechCap is not always helpful for users to express their emotions and thoughts.
From a receiver’s perspective, the majority of the participants (13/14) agreed that they could perceive the emotions of others by seeing and playing with the impact captions in VR (Q8, M=5.43, SD=0.65). Most participants (12/14) agreed that impact captions improved their understanding of others in VR (Q9, M=5.36, SD=0.74). In addition, ten participants agreed that impact captions were useful in setting the tone of conversation (Q10, M=5.00, SD=0.96). Overall, SpeechCap seems more helpful for information receivers as they felt the conversation clear and understandable, while around half of information senders thought the system failed to improve their expressions.
@@ 7.4. Results and Findings of Semi-structured Interviews
During the study, All participants completed both scenarios, playing the roles of both sender and receiver in conversations. Creative usages of impact captions in VR emerged from their play-through. In the semi-structured interviews, we discussed questions regarding the impact captions’ design space and user experiences of SpeechCap with the participants based on their individual performance. Overall, participants responded positively, describing the impact captions as “interesting and enjoyable” (P1, P2, P4, P10, P13).
Two authors analyzed the transcribed scripts of the interviews by iterative open coding  (Corbin and Strauss, 2014) and finally revealed the following findings showing the meaningfulness of our impact captions’ design space, and the usefulness of SpeechCap system.
@@ 7.4.1. F1: Textual Elements Can Effectively Support Conversations in Social VR
Participants generally agreed that visible impact captions enhanced their conversational experience in VR (P2, P3, P4, P6, P7, P11, P12, P13, P14). On one hand, text elements help users accurately express complex and profound meanings. P12 emphasized the expressive capability of impact captions: “The texts within impact captions have a high potential in expressiveness, because texts can convey complex, abstract concepts as well as accurately represent ironic or literary content, which non-textual elements cannot achieve.
@@
7.4.2. F2: Non-textual Elements are Engaging but Ambiguous
@@
In summary, the ambiguity of non-textual elements could reduce the clarity in conversations, while it is still possible to deal with the unavoidable ambiguity with design considerations, such as using the combination of multiple dimensions to complement each other.
@@ 7.4.4. F4: Needs for Identifying Keywords in Conversations
Other drawbacks of SpeechCap were exposed by the user study, too. One of the most commonly asked questions by the participants when they started using SpeechCap was about how we decided the words to be rendered as impact captions, as they usually found there were too many impact captions emerging when speaking continuously. After knowing our simple POS-tag-based filtering method to remove functional words, most participants suggested an improvement on it. P2 suggested, “Don not turn every word into an impact caption. Please extract the key information (of the speech) and only show impact captions of the relevant keywords.” Other participants, like P11, thought too much text content increased the reading and comprehension costs and then reduced the overall experience of having a conversation with SpeechCap. P11 complained, “No one wants to read that much text when using VR.” Particularly, in the multi-person conversation scenario, as more impact captions were spawned and accumulated in limited space, some participants (P1, P2) reported that they felt the field of vision was partially obstructed.
@@ 8. Discussion
8.1. Designing Interactive Visual Rhetoric for Interpersonal Communication in Social VR
As suggested by previous research, mediating interpersonal communication in virtual spaces involves supporting people in building cognitive and emotional connections  (McVeigh-Schultz and Isbister, 2021; Palmer, 1995) and exploring self-identity beyond the real world  (Freeman and Maloney, 2021; Sykownik et al., 2022; Maloney et al., 2020). However, most of the known communication methods in social VR are still imperfect  (Wei et al., 2024; Dzardanova et al., 2022; Akselrad et al., 2023; Sykownik et al., 2023; Tanenbaum et al., 2020; Baker et al., 2021).
@@ 8.1.1. Highlighting the Visual Rhetoric and Interactivity
To fill this gap, our work incorporates speech-driven captioning mechanisms into social VR with improvements on captions’ visual design and interactivity. Inspired by the original impact captions used in entertaining TV shows  (Sasamoto, 2014), we designed and implemented a visual rhetoric that integrates textual and non-textual elements to convey verbal and non-verbal information simultaneously, while enabling various human-caption interactions to support speech conversations in social VR
The user study revealed that our impact-caption-mediated approach have the potential to effectively support verbal communication in social VR (subsubsection 7.4.1). The rich interactivity provided by impact captions fosters more engaging and dynamic interpersonal interactions (subsubsection 7.4.3). However, it also faces challenges like the ambiguous non-verbal elements (subsubsection 7.4.2) and the difficulties on keyword extraction (subsubsection 7.4.4). Overall, SpeechCap provides a concrete example of how interactive and expressive visual objects can mediate communication and enhance social interactions in social VR.
@@ 8.1.2. Considering Generalizability
As for the generalizability of our impact-caption-mediated method, the aforementioned application scenarios (section 6) demonstrate that SpeechCap have the potential to be widely applied in different contexts. It can enhance emotional expressions to create engaging and intimate atmospheres, support live presentations for teaching scientific knowledge in educational scenarios, and further provide a promising captioning mechanism that can bring immersive social VR experiences to deaf and hard-of-hearing (DHH) individuals.
8.1.3. Future Directions
@@ 8.2. Understanding the Ambiguity in SpeechCap
Ambiguity in computer-mediated communication (CMC) systems is unavoidable, arising from socio-technical factors such as cultural differences, personal preferences, and technical limitations  (Stacey and Eckert, 2003; Aoki and Woodruff, 2005). Impact captions in SpeechCap also face risks of miscommunication due to inherent ambiguity in the design space.
@@
However, inevitable ambiguity in digital media design can be a double-edged sword, presenting both challenges and opportunities. Although it may introduce risks of miscommunication, ambiguity can foster creativity  (Stacey and Eckert, 2003), support negotiation  (Gaver et al., 2003), and even enhance the human-like qualities of AI agents in human-AI communication  (Liu et al., 2024). Therefore, a trade-off arises: for CMC tools, designers must balance the clarity of communication while maintaining certain ambiguity to meet other design goals  (Stacey and Eckert, 2003; Aoki and Woodruff, 2005). This trade-off leads to two directions of dealing with the ambiguity: mitigating and utilizing.
@@ 8.3.1. Presenting Verbal Information Selectively
Verbal information is fundamental in interpersonal communication  (Liu et al., 2023). However, when enhancing verbal information in social VR, it is essential to focus on selective presentation rather than displaying every spoken word. In our user study, participants suggested avoiding presenting every word in speech, since too many insignificant words could be distracting and cause unexpected visual clutter (subsubsection 7.4.4). Although we have already designed a mechanism to softly control the number of impact captions, trying to minimize the possibility of visual clutter (subsubsection 5.3.4), the POS-tag-based filtering mechanism in the text processor of SpeechCap is too simple to be an effective keyword extraction method. With the fact evolution of AI and NLP technologies, future work can explore the ways to involve advanced NLP technologies such as Large Language Models (LLMs)  (Maragheh et al., 2023) to achieve keyword extraction with accurate understandings of the speech content and the user’s intentions.
@@ 8.3.2. Using Visual Rhetoric to Present Non-verbal Information
Non-verbal information plays an important role in communication, social interactions  (McVeigh-Schultz and Isbister, 2022; Maloney et al., 2020; Aburumman et al., 2022; Liebman and Gergle, 2016), and self-presentations  (Freeman and Maloney, 2021; Sykownik et al., 2022; Zhang et al., 2022) in social VR. Our work demonstrates that rich visual elements can effectively convey non-verbal information in social VR, highlighting the value of visual rhetoric as a design strategy for communication tools. However, we currently only use simple mapping rules to decide how captions correspond to specific visual elements in SpeechCap. These rules are inadequate for sufficiently and accurately expressing the various types of non-verbal information, as pointed out by participants in the user study. Therefore, future research can start from seeking in-depth understandings of the associations between specific visual elements and non-verbal information (such as color-emotion association  (Hanada, 2018; Wilms and Oberfeld, 2018)), and create novel visual rhetoric designs with effectiveness and elegance for supporting various applications of social VR.
@@ 8.4.3. Simplifying the Hardware Requirements
8.5. Limitations and Future Work
Although SpeechCap demonstrates the potential of impact captions to enhance interpersonal communication in social VR, several limitations remain. First, the current rule-based method for determining visual appearances of captions is simplistic and may lead to ambiguity, as users interpret visual elements like colors and emojis differently. Future work could incorporate advanced NLP techniques, such as LLMs, to extract contextually meaningful keywords and adapt visual designs dynamically to user preferences. Second, the system’s reliance on multiple hardware components (e.g., laptops, external microphones, and VR headsets) limits scalability and real-world deployment. Future improvements could focus on integrating all functionalities into standalone VR headsets to simplify the setup. Finally, while the study explored dyad and group conversation scenarios, further research is needed to evaluate SpeechCap in diverse social VR applications, such as education, live streaming, and accessibility for deaf and hard-of-hearing users. Long-term studies are also required to investigate the impact of using impact captions on communication behaviors and social dynamics in virtual environments.
@@ 9. Conclusion
This research highlights the potential of impact captions as an innovative medium for enhancing interpersonal communication in social VR. By seamlessly integrating verbal and non-verbal cues through dynamic and interactive typographic elements, impact captions transform real-time conversations into playful and engaging experiences. Our proof-of-concept system, SpeechCap, demonstrates the practicality and effectiveness of this approach, while also revealing opportunities for improvement. Specifically, future work could focus on reducing ambiguity in visual designs, improving adaptability for diverse user preferences, and expanding accessibility features to ensure inclusiveness for a wide range of social VR users.